OpenStudy (anonymous):

Solve the system. y = 2 - x and y = 4x + 7 A. (1, 1) B. (1, 3) C. (-1, 3) D. (-1, 1) E. no solution

jhonyy9 (jhonyy9):

C.

jhonyy9 (jhonyy9):

y=2-x y=4x+7 ----------- 0=2-7-x-4x 0=-5-5x 5x=-5 x=-5/5 x=-1 y=2-x y=2-(-1) y=2+1 y=3 (-1,3) so C.

OpenStudy (precal):

2-x=4x+7 set up this way and solve for x jhonyy9 did you set it up this way to begin with?